# Hierarchical Reinforcement Learning for Swarm Confrontation with High Uncertainty

### v1

@@ IV-B Lower Layer for Path Planning
    Decentralized Partially Observable Markov Decision Process. Based on the allocation results of the upper layer, each pursuer needs to plan a route to chase the assigned evader and avoid collisions. We model this process as a Dec–POMDP, including the state space $S$, observation space $Z$, action space $A$, reward $R$, and state transition $P$:
@@
    State. The global state space includes information about all pursuers and evaders. It can be described as
@@
    Observation. The local observation of a pursuer contains the information of the allocated evader, its nearest neighbor and obstacle. It yields that
@@
    Action. The action consists of the velocity increment $\lVert\delta v\rVert$ and yaw $\psi$ of each UAV, which is denoted as
    $a=\left(\lVert\delta v\rVert,\psi\right).$ (10)
@@
    $r_{int}=\left\{\begin{matrix}\displaystyle-\frac{\lVert p_{i}-p_{j}\rVert}{\rho_{c,i}}+r_{a}&\lVert p_{i}-p_{j}\rVert<\rho_{c,i}\\ \displaystyle-\frac{\lVert p_{i}-p_{j}\rVert}{\rho_{c,i}}&\text{otherwise}\\ \end{matrix}\right..$ (11)
@@
    $r_{avo}=r_{avo,1}+r_{avo,2}.$ (12)
    If pursuer $i$ enters the threat zone of its nearest obstacle ($\lVert p_{i}-p_{k}\rVert<\rho_{u}+\rho_{o}+d_{thr}$), there is
    $r_{avo,1}=\frac{\lVert p_{i}-p_{k}\rVert-\left(\rho_{u}+\rho_{o}+d_{thr}\right)}{\left(\rho_{u}+\rho_{o}+d_{thr}\right)}-r_{b},$ (13)
    otherwise, $r_{avo,1}=0$. And if pursuer $i$ enters the threat zone of its nearest neighbor ($2\rho_{u}<\lVert p_{i}-p_{i}^{n}\rVert<2\rho_{u}+d_{thr}$), there is
    $r_{avo,2}=\frac{\lVert p_{i}-p_{i}^{n}\rVert-\left(2\rho_{u}+d_{thr}\right)}{\left(2\rho_{u}+d_{thr}\right)}-r_{c},$ (14)
    otherwise, $r_{avo,2}=0$. $r_{a},r_{b},r_{c}$ are constant rewards. We set a threat distance $d_{thr}$ to avoid collisions between pursuer $i$ with its neighbors and obstacles. The total reward for path planning is a linear operation consisting of the intrinsic and avoidance rewards, which is calculated as
    $r_{path}=\omega_{1}r_{int}+\omega_{2}r_{avo}.$ (15)
@@
    $p_{x,t+1}=p_{x,t}+\delta t\lVert\delta v\rVert\cos\psi,$
    $p_{y,t+1}=p_{y,t}+\delta t\lVert\delta v\rVert\sin\psi.$ (16)
@@
    $\begin{array}[]{c}L_{\theta}^{i}=\mathbb{E}_{\left(z_{t},a_{t}\right)\sim\mathfrak{B}}\lVert Q_{\theta}^{i}\left(z_{t},a_{t}\right)-y^{i}\rVert^{2},\end{array}$ (17)
    where $y^{i}=R^{i}\left(z_{t},a_{t}\right)+\gamma\max Q_{\theta^{-}}^{i}\left(z_{t+1},a_{t+1}\right)$. In addition, each agent holds a policy network $\pi_{\beta}^{i}$ which has the following loss function:
    $\begin{array}[]{c}L_{\beta}^{i}=-\mathbb{E}_{\left(z_{t},a_{t}\right)\sim\mathfrak{D}}\left[Q_{\theta}^{i}\left(z_{t},a_{t}\right)\right].\end{array}$ (18)
    Based on MADDPG, the pursuers maximize the cumulative rewards in (15) by collaborating while autonomous path planning.
@@ IV-C Hierarchical Network Interaction Method
where $\hat{s}_{t+H},\hat{r}_{t}$ denote the predicted value of $s_{t+H}$ and $r_{t}$ with the environment model, respectively. We use an ensemble neural network $(m_{\phi}^{1},...,m_{\phi}^{B})$ proposed in [29], which can be used to quantify the epistemic uncertainty and aleatoric uncertainty in the environment. Epistemic uncertainty results from the lack of sufficient training data and aleatoric uncertainty refers to the unknown enemies’ strategies and moving obstacles in this study. It takes the state–action pair as input and outputs Gaussian distribution $\mathcal{N}$ of the next state and reward. The model $m_{\phi}^{b}$ can be expressed in the following form:
@@
$\displaystyle L_{\phi}^{b}=\mathbb{E}_{\left(s_{t},a_{t}\right)\sim\mathcal{D}}\frac{1}{\lVert\mathcal{D}\rVert}\cdot\left[\frac{\log\sigma_{m_{\psi}^{b}}}{2}+\frac{\left(d_{t}-\mu_{m_{\psi}^{b}}\right)^{2}}{2\sigma_{m_{\psi}^{b}}}\right],$ (21)
@@
$H\left(s_{t},a_{t}\right)=\lfloor-\omega_{5}\hat{V}\left(s_{t},a_{t}\right)+H_{base}\rfloor,$ (24)
where $\omega_{5}$ is the weighting factor and $H\left(s_{t},a_{t}\right)$ is an integer limited in $\left[H_{min},H_{max}\right]$. Subscripts $base$, $min$, and $max$ are the settled based, minimum, and maximum values, respectively. $\lfloor x\rfloor=\max\left\{m\in\mathbb{Z}|m\leqslant x\right\}$, where $x\in\mathbb{R}$ and $\mathbb{Z}$ denotes the set of integers. In [40], it is mentioned that aleatoric uncertainty cannot be reduced with training, but we quantify the uncertainty to adjust the interaction frequency adaptively. When uncertainty $\hat{V}\left(s_{t},a_{t}\right)$ is higher, $H$ is smaller, indicating the need for more frequent target allocation by the upper layer. As the model $\mathcal{M}_{\phi}$ fits the environment, the epistemic uncertainty decreases, as well as the frequency of target allocation.
In addition, we feed the cumulative rewards to the upper after the lower executes $H$ time steps, as shown in Fig. 2. The upper layer linearly weights the cumulative rewards in (15) and the static allocation reward in (5) as the final reward, which can be denoted as
$r_{total}=\omega_{3}r_{allo}+\omega_{4}r_{path}.$ (25)
However, samples are greatly reduced since the effect of a single allocation in the upper layer can only be obtained from the feedback of the lower layer planning after $H$ steps. Here, we employ an improved model–based value expansion (IMVE) method to improve the sample utilization of the upper layer. The method performs $N$–step value estimation based on the real environment state and allows $Q_{\theta}$ of $N$ steps to converge to $Q^{*}$. Similar to $H\left(s_{t},a_{t}\right)$, $N\left(s_{t},a_{t}\right)$ is affect by the uncertainty quantified by the model $\mathcal{M}_{\phi}$. We can calculate the prospective length of $N$ in the following form:
$N\left(s_{t},a_{t}\right)=\lfloor-\omega_{6}\hat{V}\left(s_{t},a_{t}\right)+N_{base}\rfloor,$ (26)
where $N\left(s_{t},a_{t}\right)$ is an integer limited in $\left[0,N_{max}\right]$. The goal of policy in double DQN is to
@@
$\pi_{1}\left(s_{t}\right)=~{}\underset{a_{t}}{\arg\max}~{}\mathbb{E}\left[\sum_{k=t}^{\infty}{\gamma^{k}R\left(s_{k},a_{k}\right)}\right].$ (27)
We reduce the prediction interval from infinity to a visual step $N$, and replace the reward $R\left(s_{k},a_{k}\right)$ with the cumulative rewards $\sum_{i=k}^{\infty}{\gamma^{i}r_{i}}$ in (27). The strategy of IMVE can be described in the following form:
$\begin{array}[]{c}\pi_{2}\left(s_{t}\right)=\underset{a_{t:t+N}}{\arg\max}~{}\mathbb{E}\left[\sum\limits_{k=t}^{t+N}{\gamma^{k}\left(\sum\limits_{i=k}^{\infty}{\gamma^{i}r_{i}}\right)}\right].\end{array}$ (28)
It produces a local optimal solution by performing predictive control based on the single moment $t$. Through rolling iterations at different times, it generates multiple local optimal solutions and calculates the optimal value. The use of cumulative rewards compensates for the lack of the terminal reward. IMVE improves the sample utilization and training efficiency by performing $N$–step value estimation and convergence, the loss function in (6) is improved as follows:
$\begin{array}[]{c}L_{\theta}=\mathbb{E}_{\left(s_{t},a_{t}\right)\sim\mathfrak{D}}\sum\limits_{n=0}^{N-1}{\hat{\omega}\lVert Q_{\theta}\left(\hat{s}_{t+H_{n}},\hat{a}_{t+H_{n}}\right)-y_{t+H_{n}}\rVert^{2}},\end{array}$ (29)
@@
$\hat{\omega}\left(\hat{s}_{t},\hat{a}_{t}\right)=-\omega_{7}\hat{V}\left(\hat{s}_{t},\hat{a}_{t}\right)+\hat{\omega}_{base},$ (32)
where $\hat{\omega}\left(\hat{s}_{t},\hat{a}_{t}\right)$ is limited in $\left[\hat{\omega}_{min},1\right]$. By constructing the environment model $M_{\phi}$, we can quantify the epistemic uncertainty and aleatoric uncertainty, thus allowing $H$, $N$, and $\hat{\omega}$ to adjust adaptively. Moreover, it allows us to improve the sample utilization of the upper layer based on IMVE. The framework of our method is shown in Fig. 3.
@@ IV-D Integration Training Method
3:        $r_{u}\leftarrow$ Equ.(5); $\triangleright$ static allocation reward
4:        $L_{u}\leftarrow$ Equ.(6);
@@
10:        $r_{l}\leftarrow$ Equ.(15); $\triangleright$ path planning reward
11:        $L_{l}\leftarrow$ Equ.(17)–(18);
@@
16:     $t_{0}\leftarrow 1$, $H\leftarrow$ Equ.(25) $\triangleright$ initialize $H$
@@
18:        $r_{l}\leftarrow$ Equ.(15);
19:        $L_{l}\leftarrow$ Equ.(17)–(18);
20:        if $t=t_{0}+H$ then
21:           $r_{u}\leftarrow$ Equ.(26); $\triangleright$ total reward
22:           $L_{u}\leftarrow$ Equ.(6);
23:           Refresh networks $Q_{\theta_{u}}$ and $\mathcal{M}_{\phi}$;
24:           $t_{0}\leftarrow t$, $H\leftarrow$ Equ.(25) $\triangleright$ refresh $H$
25:        end if
26:        Refresh networks $Q^{i}_{\theta_{l}},\pi^{i}_{\beta_{l}}$;
@@
At the beginning of the ITM, we construct the pre–training phase for the upper layer and lower layer for $E_{p,u}$ and $E_{p,l}$ episodes, respectively. This phase allows the upper layer to learn a static allocation strategy and the lower layer to train an initial path planning policy, which supports the subsequent training phase. After the pre–training phase, we cross–train the two–layer networks for $E_{c}$ episodes. Since the value of $H$ keeps changing according to the state $s$ and allocation results $a$, we have to keep it refreshed. Also, the model $M_{\phi}$ is keeping updating at the same time in this phase.
@@ V Experiments
In this section, extensive experiments are conducted to evaluate the performance of our HRL method for solving the swarm confrontation problem in different sizes. In comparison experiments, we adopt various baselines including the expert system, game theory, heuristic, and traditional DRL algorithms. We verify the influence of adaptive frequency approach and ITM in ablation studies. Moreover, we apply the model trained under a small size to solve larger ones to investigate the generalization of our method.
@@
Figure 4: Learning curves of the pre–training for upper layer and lower layer.

### v2

@@ IV-B Lower Layer for Path Planning
    Decentralized Partially Observable Markov Decision Process. Based on the allocation result of the upper layer, each pursuer needs to plan a route to chase the assigned evader and avoid collisions. We model this process as a Dec–POMDP, including the state space $S$, observation space $Z$, action space $A$, reward $R$, and state transition $P$:
@@
    State. The global state space includes information on all pursuers and evaders. It can be described as
@@
    Observation. Considering the fast–moving characteristics of agents, instead of setting fixed distance thresholds to construct an observation vector, it is assumed that agents can observe the nearest neighbor. The local observation of a pursuer contains information on the allocated evader, its nearest neighbor, and obstacle. It yields that
@@
    Action. The action consists of the velocity magnitude $\lVert v\rVert\in[0,\lVert v\rVert_{\max}]$ and velocity direction $\psi\in[-\pi,\pi]$ of pursuer $i$, which is denoted as
    $a=\left(\lVert v_{i}\rVert,\psi_{i}\right).$ (15)
@@
    $r_{\text{int}}=\left\{\begin{matrix}\displaystyle-\frac{\lVert p_{i}-p_{j}\rVert}{\rho_{c,i}}+r_{a}&\lVert p_{i}-p_{j}\rVert<\rho_{c,i}\\ \displaystyle-\frac{\lVert p_{i}-p_{j}\rVert}{\rho_{c,i}}&\text{otherwise}\\ \end{matrix}\right..$ (16)
@@
    $r_{\text{avo}}=r_{\text{avo},1}+r_{\text{avo},2}.$ (17)
    If pursuer $i$ enters the threat zone of its nearest obstacle $k$ ($\lVert p_{i}-p_{k}\rVert<\rho_{u}+\rho_{o}+d_{\text{thr}}$), there is
    $r_{\text{avo},1}=\frac{\lVert p_{i}-p_{k}\rVert-\left(\rho_{u}+\rho_{o}+d_{\text{thr}}\right)}{\left(\rho_{u}+\rho_{o}+d_{\text{thr}}\right)}-r_{b},$ (18)
    otherwise, $r_{\text{avo},1}=0$. And if pursuer $i$ enters the threat zone of its nearest neighbor ($2\rho_{u}<\lVert p_{i}-p_{i}^{n}\rVert<2\rho_{u}+d_{\text{thr}}$), there is
    $r_{\text{avo},2}=\frac{\lVert p_{i}-p_{i}^{n}\rVert-\left(2\rho_{u}+d_{\text{thr}}\right)}{\left(2\rho_{u}+d_{\text{thr}}\right)}-r_{c},$ (19)
    otherwise, $r_{\text{avo},2}=0$. $r_{a},r_{b},r_{c}$ are constant rewards. We set a threat distance $d_{\text{thr}}$ to avoid collisions between pursuer $i$ with its neighbors and obstacles. The final reward for path planning is a linear operation consisting of the intrinsic reward and avoidance reward, which can be calculated as
    $r_{\text{path}}=\omega_{2}r_{\text{int}}+\left(1-\omega_{2}\right)r_{\text{avo}}.$ (20)
    where $\omega_{2}$ is the weighting factor between the intrinsic reward and avoidance reward.
@@
    $p_{x,t+1}=p_{x,t}+\delta t\lVert v\rVert\cos\psi,$
    $p_{y,t+1}=p_{y,t}+\delta t\lVert v\rVert\sin\psi,$ (21)
    where $\delta t$ is the length of time step.
@@
    $\begin{array}[]{c}L_{\theta}^{i}=\mathbb{E}_{\left(z_{t},a_{t}\right)\sim\mathcal{D}}\lVert Q_{\theta}^{i}\left(z_{t},a_{t}\right)-y^{i}\rVert^{2},\end{array}$ (22)
    where $y^{i}=R^{i}\left(z_{t},a_{t}\right)+\gamma\max Q_{\theta^{-}}^{i}\left(z_{t+1},a_{t+1}\right)$. In addition, each agent holds a policy network $\pi_{\beta}^{i}\left(z^{i}_{t}\right)$ which has the following loss function:
    $\begin{array}[]{c}L_{\beta}^{i}=-\mathbb{E}_{\left(z_{t},a_{t}\right)\sim\mathcal{D}}\left[Q_{\theta}^{i}\left(z_{t},a_{t}\right)\right].\end{array}$ (23)
    The subscript $\beta$ is the weighting factor of the policy network. Based on MADDPG, the pursuers maximize the cumulative rewards in (20) by collaborating while autonomous path planning.
@@ IV-C Hierarchical Network Interaction Method
where $\hat{s}_{t+H},\hat{r}_{t}$ denote the predicted value of $s_{t+H}$ and $r_{t}$ with the environment model, respectively. We use an ensemble neural network $(m_{\phi}^{1},...,m_{\phi}^{B})$ proposed in [29], which can be used to quantify the epistemic and aleatoric uncertainty in the environment. Epistemic uncertainty results from the lack of sufficient training in the lower layer and aleatoric uncertainty refers to the unknown enemies’ strategies and moving obstacles in this study. It takes the state–action pair as input and outputs Gaussian distribution $\mathcal{N}$ of the next state and reward. The model $m_{\phi}^{b}$ can be expressed in the following form:
@@
$\displaystyle L_{\phi}^{b}=\mathbb{E}_{\left(s_{t},a_{t}\right)\sim\mathcal{D}}\frac{1}{\lVert\mathcal{D}\rVert}\left[\frac{\log\sigma_{m_{\psi}^{b}}}{2}+\frac{\left(d_{t}-\mu_{m_{\psi}^{b}}\right)^{2}}{2\sigma_{m_{\psi}^{b}}}\right],$ (26)
@@
$H\left(s_{t},a_{t}\right)=\lfloor-\omega_{3}\hat{V}\left(s_{t},a_{t}\right)+H_{\text{base}}\rfloor,$ (29)
where $\omega_{3}$ is a weighting factor and $H\left(s_{t},a_{t}\right)$ is an integer limited in $\left[H_{\min},H_{\max}\right]$. Subscripts base, $\min$, and $\max$ are the settled based, minimum, and maximum values, respectively. $\lfloor x\rfloor=\max\left\{m\in\mathbb{Z}|m\leqslant x\right\}$, where $x\in\mathbb{R}$ and $\mathbb{Z}$ denotes the set of integers. In [40], it is mentioned that aleatoric uncertainty cannot be reduced with training, but we quantify the uncertainty to adjust the interaction frequency adaptively. When uncertainty $\hat{V}\left(s_{t},a_{t}\right)$ is higher, $H$ is smaller, indicating the need for more frequent target allocation by the upper layer. As the model $\mathcal{M}_{\phi}$ fits the environment, the epistemic uncertainty in the lower layer decreases, as well as the frequency of target allocation.
In addition, we feed the cumulative rewards to the upper after the lower executes $H$ time steps, as shown in Fig. 2. The static allocation reward in (10) and the cumulative path rewards in (20) encourage agents to allocate reasonable targets and plan feasible paths, respectively. Therefore, the upper layer linearly weights the above rewards as the total reward for target allocation, which can be denoted as
$r_{\text{total}}=H\cdot r_{\text{allo}}+\sum_{1}^{H}{r_{\text{path}}}+r_{\text{capt}},$ (30)
where $r_{\text{capt}}$ is an additional reward equal to the number of captured evaders during these $H$ time steps. Based on the total reward, the upper layer is updated by making a trade–off among the above three objective rewards simultaneously. However, samples are greatly reduced since the effect of a single allocation in the upper layer can only be obtained from the feedback of the lower layer planning after $H$ steps. Here, we employ an improved model–based value expansion (IMVE) method to improve the sample utilization of the upper layer. The method performs $N$–step value estimation based on the real environment state and allows $Q_{\theta}$ of $N$ steps to converge to $Q^{*}$. Similar to $H\left(s_{t},a_{t}\right)$, $N\left(s_{t},a_{t}\right)$ is affect by the uncertainty quantified by the model $\mathcal{M}_{\phi}$. We can calculate the prospective length of $N$ in the following form:
$N\left(s_{t},a_{t}\right)=\lfloor-\omega_{4}\hat{V}\left(s_{t},a_{t}\right)+N_{\text{base}}\rfloor,$ (31)
where $\omega_{4}$ is a weighting factor and $N\left(s_{t},a_{t}\right)$ is an integer limited in $\left[1,N_{\max}\right]$. The goal of policy in double DQN is to
@@
$\pi_{1}\left(s_{t}\right)=~{}\underset{a_{t}}{\arg\max}~{}\mathbb{E}\left[\sum_{m=t}^{\infty}{\gamma^{m}R\left(s_{m},a_{m}\right)}\right],$ (32)
[…]
$\begin{array}[]{c}\pi_{2}\left(s_{t}\right)=\underset{a_{t:t+N}}{\arg\max}~{}\mathbb{E}\left[\sum\limits_{m=t}^{t+N}{\gamma^{m}\left(\sum\limits_{g=m}^{\infty}{\gamma^{g}r_{g}}\right)}\right].\end{array}$ (33)
It produces a locally optimal solution by performing predictive control based on the single moment $t$. Through rolling iterations at different times, it generates multiple locally optimal solutions and calculates the optimal value. The use of cumulative rewards compensates for the lack of the terminal reward. IMVE improves the sample utilization and training efficiency by performing $N$–step value estimation and convergence, the loss function in (11) is improved as follows:
$\begin{array}[]{c}L_{\theta}=\mathbb{E}_{\left(s_{t},a_{t}\right)\sim\mathcal{D}}\sum\limits_{n=0}^{N-1}{\hat{\omega}\lVert Q_{\theta}\left(\hat{s}_{t+H_{n}},\hat{a}_{t+H_{n}}\right)-y_{t+H_{n}}\rVert^{2}},\end{array}$ (34)
@@
$\hat{\omega}\left(\hat{s}_{t},\hat{a}_{t}\right)=-\omega_{5}\hat{V}\left(\hat{s}_{t},\hat{a}_{t}\right)+\hat{\omega}_{\text{base}},$ (37)
where $\omega_{4}$ is a weighting factor and $\hat{\omega}\left(\hat{s}_{t},\hat{a}_{t}\right)$ is limited in $\left[\hat{\omega}_{\min},1\right]$. By constructing the environment model $M_{\phi}$, we can quantify the epistemic and aleatoric uncertainty, thus allowing $H$, $N$, and $\hat{\omega}$ to adjust adaptively. Moreover, it allows us to improve the sample utilization of the upper layer based on IMVE. The framework of our method is shown in Fig. 3.
@@ IV-D Integration Training Method
3:        $r_{u}\leftarrow$ Eq. (10); $\triangleright$ static allocation reward
4:        $L_{u}\leftarrow$ Eq. (11);
@@
10:        $r_{l}\leftarrow$ Eq. (20); $\triangleright$ path planning reward
11:        $L_{l}\leftarrow$ Eq. (22)–(23);
@@
16:     $t_{0}\leftarrow 1$, $H\leftarrow$ Eq. (29) $\triangleright$ initialize $H$
@@
18:        $r_{l}\leftarrow$ Eq. (20);
19:        $L_{l}\leftarrow$ Eq. (22)–(23);
20:        Refresh networks $Q^{i}_{\theta_{l}},\pi^{i}_{\beta_{l}}$;
21:        if $t=t_{0}+H$ then
22:           $r_{u}\leftarrow$ Eq. (30); $\triangleright$ total reward
23:           $L_{u}\leftarrow$ Eq. (11);
24:           Refresh networks $Q_{\theta_{u}}$ and $\mathcal{M}_{\phi}$;
25:           $t_{0}\leftarrow t$, $H\leftarrow$ Eq. (29) $\triangleright$ refresh $H$
26:        end if
@@
At the beginning of the ITM, we construct the pre–training phase for the upper layer and lower layer for $E_{p,u}$ and $E_{p,l}$ episodes, respectively. This phase allows the upper layer to learn a static allocation strategy and the lower layer to train an initial path planning policy, which supports the subsequent training phase. After the pre–training phase, we cross–train the two–layer networks for $E_{c}$ episodes. Since the value of $H$ keeps changing according to the state $s$ and allocation result $a$, we have to keep it refreshed. Also, the model $M_{\phi}$ is keeping updating at the same time in this phase.
@@ V Experiments
In this section, extensive experiments are conducted to evaluate the performance of our HRL method for solving the swarm confrontation problem in different sizes. In comparison experiments, we adopt various baselines including the expert system, game theory, heuristic, and traditional DRL algorithms. We verify the influence of the adaptive frequency approach and ITM in ablation studies. Moreover, we apply the model trained under a small size to solve larger ones to investigate the generalization of our method. Finally, we deploy the model after integration training in simulations to the real–robot system.
@@
Figure 4: Learning curves of the pre–training for upper layer and lower layer. (a) Target allocation. (b) Path planning.